Please can you help solve this problem for me. The point (x,y) is equidistant from (4,-1) and (-2,3). Express the answer in algebraic equation
Found 2 solutions by LinnW, Alan3354:
Answer by LinnW(1048)

We are looking for the midpoint between the two points.
For points (x1,y1) and (x2,y2) the midpoint is
( (x1+x2)/2 , (y1+y2)/2 )
For our point values (4,-1) and (-2,3)
this becomes
( (4+(-2))/2 , (-1+3)/2 )
( 2/2 , 2/2 )
( 1 , 1 )

Answer by Alan3354(69443)

Please can you help solve this problem for me. The point (x,y) is equidistant from (4,-1) and (-2,3). Express the answer in algebraic equation
---------------
There are an infinite # of points. It's the perpendicular bisector thru (1,1)
---
Find the slope of the line thru the given points.
m = diffy/diffx = 4/-6 = -2/3
The slope of lines perpendicular is the negative inverse, 3/2
-----
Find an equation of the line thru (1,1) with a slope of 3/2
Use y = mx+b and the point to find b
---
1 = (3/2)*1 + b
b = -1/2
--> y = 3x/2 - 1/2